Dublin University Rifle Club

Version 1.1 Date : May 16, 2001

DURC Competition Attendance Report

Air Rifle Nationals 2001

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue UCD Rifle Club, Belfield
Date May 12/13, 2001
Discipline 10m Air Rifle

DURC Attendees

Mark Dennehy Ben Jones

Keith White

Scores

Mark Dennehy 513
Ben Jones 505

488

Prizes Won

Ben Jones 3^{rd} Class D

Report

The 2001 Air Rifle National Championships were held in the UCD rifle range on May 12^{th} and 13^{th}. DURC were asked to assist with the logistics for the competition by the National Target Shooting Association on the basis of the standard set by the DURC Air Rifle Open earlier this season - an excellent compliment given that our previous Open was deregistered due to the poor facilities in Luce Hall.

Competing were Mark Dennehy and Ben Jones; Keith White was the offical scorer and statistician for the competition. Our computer was transported to the UCD range on the Wednesday prior to the competition to allow for setting up and debugging. Unfortunately we were unable to secure an additional laptop and projector as we had for our Open, so a second monitor was connected to our PC to allow one screen for editing of results and one for display. This was accomplished using Keith White’s SCORP program - we expect this program to be adopted by the rest of the country once Keith releases his final version, given the interest expressed by clubs from Rathdrum to Comber to Wales.

The competition consisted of an Open, which comprised two 60-shot matches, and the Nationals itself, which was a single 60-shot match, shot concurrently with the Open. Mark shot a 513 (ex600) on the first day but due to an admin foul-up was disqualified from the Open when his second match had only 55 shots, one of his target cards being lost in the meleé that Sundays shooting proved to be. Ben Jones shot very well, scoring a 505 on his first day, which is the first time he has broken 500 - a major milestone in Air Rifle. Sadly neither of our competitors made the ten-shot finals due to the excellent standards of the other shooters - but in their defence, they were competing against two Olympic shooters, all but two of the Irish National Squad, and two of the Welsh Squad as well! Ben Jones did manage to get third place in Class D however, beating off virtually all of the Wilkinstown club to do so (some twenty shooters!); and he has now been promoted to Class C; and since 510 is the cutoff for Class B, he may well be promoted to Class B within two more Air Rifle competitions. This marks the end of the 2000-2001 Air Rifle season: The next Air Rifle shoot in Ireland is the East Antrim Festival in early October, which we plan to attend.

Misé le meas,

Mark Dennehy, DURC Secretary